OpenStudy (anonymous):

why is this wrong? a)185e^(ln(140/185)/30*45)? A roasted turkey is taken from an oven when its temperature has reached 185 Fahrenheit and is placed on a table in a room where the temperature is 75 Fahrenheit. (a) If the temperature of the turkey is 140 Fahrenheit after half an hour, what is its temperature after 45 minutes? b) When will the trukey cool to 100 Fahrenheit?

OpenStudy (anonymous):

T(t) = Ts + (To - Ts)*e^(-kt). Where did you get your value for k? When I did the math (you can see my work), I got 0.0070767. . . With that value for k, both equations work out.

OpenStudy (anonymous):

or this...... 25 = 110 * e^(-0.0070767t) 25/110 = e^(-0.0070767t) ln(25/110) = -0.0070767t t = ln(25/110)/-0.0070767 t = 209.3623 minutes

OpenStudy (anonymous):

i just used the orginal equation, and that value, 209.3623 for part b isnt correct. i have looked the problem both ways and its still wrong